How much work is needed to decreases the distance between a +15 µC charge





and a -20 µC charge from 1 m to 0.25 m?

1
Expert's answer
2022-03-20T18:49:02-0400

The work done is equal to change in potential energy of system

"W=\Delta E_p=k\frac{q_1q_2}{r_2}-k\frac{q_1q_2}{r_1}"

"W=9*10^9*15*10^{-6}*(-20*10^{-6})\\
\times\left(\frac{1}{0.25}-\frac{1}{1.00}\right)=-8.1\:\rm J"